OpenStudy (anonymous):

solve each system of equations: 1) 5x+4y-z=1 2x-2y+z=1 -x-y+z=2 i got x=0, y=-1, z=-1, is that correct?

OpenStudy (anonymous):

For the first equation: 0 - 4 + 1 does not equal 1

OpenStudy (anonymous):

well then what should i do?

OpenStudy (anonymous):

These get messy ... start again ... which variable do you want to eliminate first?

OpenStudy (anonymous):

I suggest eliminating z: From the first two equations we get 7x + 2y = 2 and from the first and the third we get 4x + 3y = 3

OpenStudy (anonymous):

oh i got 7x-2y=2 and 4x-3y=3...

OpenStudy (anonymous):

That happens to me too .. 4y - 2y = 2y and 4y - y = 3y

OpenStudy (anonymous):

oh ok so anyways i tried it with your equations and i still got x=0...

OpenStudy (anonymous):

y=1 and z=3?

OpenStudy (anonymous):

Yes, x = 0

OpenStudy (anonymous):

Yes, the solution is (0, 1, 3)

OpenStudy (anonymous):

oh ok thanks so much!